MATEMATISKA INSTITUTIONEN STOCKHOLMS UNIVERSITET

Avd. Matematik Jörgen Backelin Kommutativ algebra HT 2003

If you want documentedly to have completed the course and its examination before the end of Januar (2004), then you should try to give me all or at least almost all your answers before the 6’th of Januar.

The second and last round consists of the following exercises from Atiyah-Macdonald: 3.12, 4.16, 6.5, 6.6, 6.8, 6.9, 7.11, 7.21, 8.2, 8.4. (Note the printing error in the hints to exercise 3.12.iv): you may have use of “Chapter 2, Exercise 15 and Exercise 20”.)